A football with a mass of 0.4 kg is thrown with a force of 60.0 N. Discounting friction, what is the acceleration of the ball?


1
Expert's answer
2021-02-06T10:05:03-0500
"a=\dfrac{F}{m}=\dfrac{60.0\ N}{0.4\ kg}=150\ \dfrac{m}{s^2}."